Consider ODE

(3x^2y^2+x^2)dx+(2x^3y+y^2)dy=0

Find genaral solution


1
Expert's answer
2021-06-29T07:51:08-0400
"(3x^2y^2+x^2)dx+(2x^3y+y^2)dy=0"

"P=3x^2y^2+x^2, \dfrac{\partial P}{\partial y}=6x^2y"

"Q=2x^3y+y^2, \dfrac{\partial Q}{\partial x}=6x^2y"

"\dfrac{\partial P}{\partial y}=6x^2y=\dfrac{\partial Q}{\partial x}"

"\dfrac{\partial u}{\partial x}=P(x, y), \dfrac{\partial u}{\partial y}=Q(x,y)"

"u(x,y)=\int(3x^2y^2+x^2)dx+\varphi(y)"

"=x^3y^2+\dfrac{1}{3}x^3+\varphi(y)"

"\dfrac{\partial u}{\partial y}=2x^3y+\varphi'(y)=2x^3y+y^2"

"\varphi'(y)=y^2"

"\varphi(y)=\dfrac{1}{3}y^3+C_1"

The general solution of the differential equation


"(3x^2y^2+x^2)dx+(2x^3y+y^2)dy=0"

is given by


"x^3y^2+\dfrac{1}{3}x^3+\dfrac{1}{3}y^3=C"